# SLED: A Speculative LLM Decoding Framework for Efficient Edge Serving

### v1

@@ V Conclusion
This paper presented the SLED framework, a novel distributed decoding solution designed for large language model deployment at the edge. Our extensive evaluation demonstrated that SLED significantly improves scalability, computational efficiency, energy efficiency, and robustness compared to traditional centralized approaches. The integration of speculative local drafting and centralized verification establishes a balanced trade-off between latency and energy consumption, making SLED particularly suitable for latency-sensitive edge applications. It’s proved in this position paper that the SLED is more than a decoding enhancement—it opens the door to more foundational and elastic approach to resource-aware LLM serving at the edge.
@@ VI Future Work
Future research directions include enhancing the adaptive capabilities of SLED for dynamic environments, such as network condition and resource aware verification strategy. Additionally, exploring the smart queue and batching strategy on the edge server for latency-sensitive task and better server utilization will be another interest topic to focus on. Lastly, extending SLED’s applicability to multi-modal and real-time streaming scenarios could further broaden its practical impact and robustness in complex edge computing landscapes.
[23]

### v2

@@ V Conclusion
This paper presented the SLED framework, a novel distributed decoding solution designed for LLM deployment at the edge. Our extensive evaluation demonstrated that SLED significantly improves system throughput, capacity and cost efficiency compared to traditional centralized approaches. The integration of speculative local drafting and centralized verification establishes a balance of computational workload, making SLED particularly suitable for bringing LLMs towards the edge of network. It’s proved in this position paper that the SLED is more than a decoding enhancement—it opens the door to more foundational and elastic approach to resource-aware LLM serving at the edge.
@@ VI Future Work
Future research directions include exploring the smart queue and batching strategy on the edge server for latency-sensitive task and better server utilization. Additionally, enhancing the adaptive capabilities of SLED for dynamic environments, such as extending SLED’s applicability to multi-modal scenarios will be another interesting topic to focus on. Lastly, network condition and resource aware verification strategy could further broaden its practical impact in complex edge computing landscapes.